An explorer is caught in a whiteout (in which the snowfall is so thick that the ground cannot be distinguished from the sky) while returning to base camp. He was supposed to travel north for 5.6 km, but when the snow clears, he discovers that he actually traveled 7.8 km at 50° east of north. How far and in what direction must he now travel to reach base camp?


1
Expert's answer
2021-05-30T13:28:28-0400
"(0,5.6)=(x,y)+(7.8\sin{50},7.8\cos{50})\\(x,y)=(-6,0.6)\ km\\s=6.0\ km\\\theta=5.7\degree\text{ W of N}"